Abbreviated income statements for Pop co. follow:

2018 2017
Sales 210,000 165,000
COGS 149,000 105,000
Operating Expense 27,500 18,000
Income Tax Expense 10,800 11,100
Net Income 22,700 30,900


Using Vertical analysis , which of the following is true?
a. 2018 operating Expenses as a percent of Sales is 4.91%
b. 2017 Net incom as a percent of sales is 4.91%
c. Both a,b are true.
d. Nethier a, nor b are true.

H2O Corporation had sales of $547,000 cost of goods sold of $348,000, and avergae inventory of $65,000. what is the company's inventory turnover ratio?
a. 0.19x
b. 3.06x
c. 5.35x
d. 8.42x

the Tea company only has common stockholders. Last year , the company earned $47,000 on sales of $163,200: the company average 18,050 outstanding common shares . On the most recent tranding , the common shares were selling at $75. What is this company's earning per share?
a. $11.65 per share.
b. $9.04 per share
c. $2.60 per share.
d. $0.38 per share.

The Orange Company reported net sales of $427,000. the company has average total assets of $928,600 of which $475,000 is property, plant, and equipment and the remainder is in current assets. What is the compan's assets tuenover ratio?

a. 0.46x

b. 0.90x

c. 0.94x

d. 2.17x

Which of the following cash flows wold included under the investing activites section of the cash flow statemnt? Assum indirect method used

a. Changes in the Accounts Recivable

b. Changes in (Gross) long-term Assets

c. Net income.

d. Cash dividend
Answer #1

Answer to 1st Question:

The correct answer is d.

Reasoning:

Vertical analysis of an income statement is done considering sales as base figure.

Operating expenses of 2018 as a % of sales = (27,500/210,000)*100 = 13.1%

Hence, the statement made in option a is incorrect.

Net Income of 2017 as a % of sales = (30,900/165,000)*100 = 18.73%

Hence, the statement made in option b is incorrect and also in option c.

Answer to 2nd Question:

Option c is correct.

Reasoning:

Inventory Ratio = Cost of goods sold/Average Inventory

Inventory Ratio = 348,000/65,000 = 5.35

Answer to 3rd question:

Option c is correct.

Reasoning:

EPS = Net Income/Average shares outstanding

EPS = 47,000/18,040 = 2.60 per share